### v1

# Unsupervised Text Embedding Space Generation Using Generative Adversarial Networks for Text Synthesis

Abstract
Generative Adversarial Networks (GAN) is a model for data synthesis, which creates plausible data through the competition of generator and discriminator. Although GAN application to image synthesis is extensively studied, it has inherent limitations to natural language generation. Because natural language is composed of discrete tokens, a generator has difficulty updating its gradient through backpropagation; therefore, most text-GAN studies generate sentences starting with a random token based on a reward system. Thus, the generators of previous studies are pre-trained in an autoregressive way before adversarial training, causing data memorization that synthesized sentences reproduce the training data. In this paper, we synthesize sentences using a framework similar to the original GAN. More specifically, we propose Text Embedding Space Generative Adversarial Networks (TESGAN) which generate continuous text embedding spaces instead of discrete tokens to solve the gradient backpropagation problem. Furthermore, TESGAN conducts unsupervised learning which does not directly refer to the text of the training data to overcome the data memorization issue. By adopting this novel method, TESGAN can synthesize new sentences, showing the potential of unsupervised learning for text synthesis. We expect to see extended research combining Large Language Models with a new perspective of viewing text as an continuous space.
Keywords:
Text Embedding Space Text Synthesis Generative Adversarial Network Adversarial Learning Data Memorization Backpropagation.
1 Introduction
Generative Adversarial Networks (GAN) proposed by Goodfellow et al. (2014) [7] is a popular model for synthesizing data. GAN aims to generate plausible data, capturing real data distribution in an unsupervised way through the competition of generator and discriminator. When GAN was first proposed, it mainly focused on image synthesis, and extensive research was dedicated to achieving quality synthetic data results [2, 22, 13, 12]. The generator of GAN learns implicit density based on the loss of the discriminator without directly referencing to the training data; therefore, GAN can avoid data memorization that the model reproduces the training data. Also, GAN can synthesize various data with random noise instead of using specific starting point such as a start token.
Applications of GAN to image synthesis have been widely studied, but it has had limitations to natural language generation. As natural language is composed of discrete tokens, the generator has an issue with updating its gradient directly through backpropagation. The backpropagation problem of text-based GAN was first discussed by Yu et al. (2017) [29] and a lot of follow-up text-GAN research tried to deal with the backpropagation problem, using gradient policy-based reinforcement learning with a reward system. Also, the previous text-GAN approaches have to pre-train the generator with supervised learning (autoregressive) before adversarial training due to the convergence issue of the generator [29]. Accordingly, we found that the generators of previous text-GAN approaches reproduce the training data (data memorization) when synthesizing text due to the pre-training and this is a major issue of the generative models.
This paper suggests a novel framework, Text Embedding Space Generative Adversarial Networks (TESGAN), to enable backpropagation and avoid data memorization. The gradient backpropagation is possible in TESGAN because the model generates continuous text embedding spaces from random noise instead of discrete tokens. In addition, TESGAN does not require the supervised pre-trained generator adopted in previous research for text synthesis. In other words, TESGAN does not directly refer to the training data because it trains in an unsupervised way similar to the original GAN framework. Thus, data memorization does not arise while synthesizing sentences. We compute Fréchet BERT Distance (FBD) and Multi-sets-Jaccard (MSJ) proposed by Alihosseini et al. (2019) [1], which evaluate both quality and diversity of synthetic text simultaneously. Furthermore, TESGAN-based models show the best Language Model score (LM) [20, 4] compared to baselines. We also calculate data memorization ratio and show the synthesized sentences to assess the potential of unsupervised learning and continuous embedding spaces for text synthesis.
2 Related Works
The most common method of sentence generation is to use an autoregressive-based language model via teacher forcing [28]. For example, there are extensive studies on models using recurrent neural networks (RNN) with Long Short-Term Memory (LSTM) cells proposed by Hochreiter and Schmidhuber (1997) [11]. Using LSTM, Graves (2013) [8] successfully generated handwriting by predicting a sequence, and Wen et al. [27] synthesized sentences under specific conditions. Bowman et al. (2016) [3] generated text after learning text embedding spaces with an autoregressive-based LSTM model and variational autoencoder (VAE) architecture [15]. Policy Gradient with BLEU (PG-BLEU) computes the BLEU [21] of synthesized sentences and takes them as a reward when updating the generator with policy gradient.
Many investigations were conducted to use GAN for text synthesis. Sequence GAN (SeqGAN) proposed by Yu et al. (2017) [29] attempted to solve the backpropagation problem by using gradient policy-based reinforcement learning with a reward system. However, since SeqGAN has reward sparsity issue, Lin et al. (2017) [18] proposed RankGAN which replaces the previous regression-based discriminator with a novel ranker. RankGAN trains the discriminator which gives a higher score to more realistic sentences. In addition, Che et al. (2017) [5] proposed Maximum Likelihood Augmented Discrete GAN (MaliGAN) that synthesizes text by minimizing Kullback-Leibler divergence [16]. TextGAN [30] proposed kernel-based moment-matching which forces empirical distributions of real and synthetic text using LSTM and Convolutional Neural Networks (CNN) for the generator and the discriminator, respectively. LeakGAN [9] used a similar architecture to TextGAN, but they alleviated sparseness and lack of intermediate information by providing leaked information from the discriminator. The aforementioned text-GAN models need the first token to synthesize text because of an autoregressive generator. TESGAN is distinct from these models in that it generates text embedding space directly from random noise without selecting tokens. Our TESGAN is the first text-GAN model which learns without an autoregressive generator.
3 Text Embedding Space GAN
TESGAN aims to generate seeds required to synthesize plausible text. The generated seeds (fake seeds) from the generator and the real seeds from the real text are passed over to the discriminators to train the GAN framework.
Figure 1: Illustration of the seed interpretation model. After pre-training the seed interpretation model using multi-turn sentences, the model is used to synthesize text after freezing the parameters. Regardless of pre-training and text synthesis, all data pass through the real seed making process (sigmoid embedding) before entering the model as input. The $[PAD]$ tokens after the $[SEP]$ tokens are omitted in the left part for easy understanding.
3.1 Seed for Text Synthesis
We denote a text sequence as $S=w_{1},...,w_{T}$ ($T$ is the sequence length). An autoregressive-based language model calculates the probability of the text sequence $S$ as a product of conditional probabilities. If we assume that $S$ is a complete sentence, then their sequences $D=S_{1},...,S_{N}$ ($N$ is the dialogue length) can be seen as multi-turn sentences. Let $S_{1}=w_{1},...,w_{m}$ and $S_{2},...,S_{N}=w_{m+1},...,w_{M}$ denote the first sentence and the subsequent sentences, respectively ($M$ is the total length of the multi-turn sentences $D$). The following sentences after the first sentence can be predicted from a product of conditional probabilities:
$\small p(S_{2},...,S_{N}|S_{1})=\prod_{i=m+1}^{M}p(w_{i}|w_{1},...,w_{i-1})$ (1)
In other words, the first sentence can generate subsequent sentences using an autoregressive-based language model trained with multi-turn sentences. Therefore, the first sentence can be used as a seed. Meanwhile, the generator of TESGAN generates the first sentence as a continuous embedding space instead of discrete tokens to enable gradient backpropagation. Consequently, a continuous embedding space of the first sentence is defined as a seed.
3.2 Seed Interpretation Model
The seed interpretation model $f_{\theta}(\cdot)$ is used to predict subsequent sentences a seed makes. Therefore, the seed interpretation model must be trained with multi-turn sentences in an autoregressive way as same as general language modeling, as shown in Figure 1. The model has been trained by data consisting of multi-turn sentences $D=S_{1},...,S_{N}$, and the maximum length of each sentence is the same as $L$, that is, the number of tokens in $D$ is $N*L$. When constructing the multi-turn sentence data, the $[CLS]$ special token is inserted only at the beginning of the first sentence, and each sentence is distinguished by adding the $[SEP]$ special token at the end. If the length of a tokenized sentence is less than $L$, the $[PAD]$ tokens are used for padding the sentence:
$\small\begin{gathered}S_{1}=w_{1}^{1},...,w_{|S_{1}|}^{1},...,w_{L}^{1}\\ (w_{1}^{1}=[CLS],w_{|S_{1}|}^{1}=[SEP],w_{|l>S_{1}|}^{1}=[PAD])\end{gathered}$ (2)
$\small\begin{gathered}S_{i(i>1)}=w_{1}^{i},...,w_{|S_{i}|}^{i},...,w_{L}^{i}\\ (w_{|S_{i}|}^{i}=[SEP],w_{|l>S_{i}|}^{i}=[PAD])\end{gathered}$ (3)
where the $S_{1}$ is seed sentence. Let $H_{real}$ and $H_{fake}$ denote the real and the fake seeds, respectively, from TESGAN. As shown in Figure 1, the real seed is an embedding space of a sentence obtained by applying the sum of the token embedding layer and the positional embedding to the sigmoid function. Its dimension is (sequence length * embedding dimension). Any sentence can be used as a seed for text synthesis as long as the seed interpretation model is trained with multi-turn sentences because every sentence can exist before any other sentence. Therefore, all sentences can be used as a seed:
$\small\begin{gathered}H_{real}=\sigma\big{(}W_{emb}(S_{1})+W_{pos}(S_{1})\big{)}\\ \approx\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\end{gathered}$ (4)
$H_{real}$ from the real text can be seen as continuous spaces similar to an image. Therefore, if the generator can create $H_{fake}$ skillfully imitating distributions of $H_{real}$, $H_{fake}$ also can generate appropriate subsequent sentences (a.k.a synthetic text). No matter how excellent $H_{fake}$ are generated by the generator, it is useless if it cannot be interpreted; therefore, training the seed interpretation model is crucial. In other words, the seed interpretation model has to predict the next sentence $S_{n+1}$ properly in multi-turn cases as Equation 5. Consequently, text synthesis is conducted as the seed passes through the seed interpretation model to predict the next sentence as shown in Figure 1.
$\small S_{n+1}=f_{\theta}\Big{(}\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\Big{)}$ (5)
GPT-2 [23] is used as the seed interpretation model and it is only used to synthesize sentences from seeds with frozen parameters. Thus the seed interpretation model never affects the training of the generator and the discriminator during adversarial training, and vice versa. More detailed specifications are explained in Appendix A.
3.3 Generator
Both real and fake seeds ($H_{real}$ and $H_{fake}$) are essential for adversarial training. The real seeds can be obtained from the real text as described in the seed interpretation model, and the fake seeds are generated by the generator. In most text-GAN models reported so far, fake sentences are obtained from a text-based pre-trained autoregressive generator. Thus, data memorization that several synthetic sentences reproduce the training data occurs. To prevent data memorization, our generator does not use a pre-trained autoregressive-based model and does not explicitly refer to the text in the training data during adversarial training. Our generator aims to create suitable fake text embedding spaces (fake seeds, $H_{fake}$) in an unsupervised way (GAN framework) by referencing continuous spaces (real seeds). The embedding space created by the generator has the same dimension of (sequence length * embedding dimension), just as the real seed.
The generator $g_{\phi}(\cdot)$ is composed of two convolutional layers and generates seeds from the uniform distribution noise $X$ with an interval of $[-10,10)$ to make diverse forms of the seeds. Also, using random noise has the advantage of not having to select the first token in the text synthesis process after model learning. The final $H_{fake}$ can be obtained by Equation 6:
$\small H_{fake}=g_{\phi}\big{(}X\sim U(-10,10)\big{)}$ (6)
Moreover, we compare one more additional model, perturbed TESGAN (P-TESGAN). P-TESGAN creates perturbed seeds by adding zero-centered normal distribution noise $z$ to the generator output. P-TESGAN is expected to learn more robustly by perturbating the generator output. Further details of the generator are explained in Appendix A.
3.4 Objective Functions
Since the generator does not refer to text while adversarial training, its performance is determined by the loss of the discriminator. Thus, we propose four types of loss to update the parameters of the generator and the discriminator.
3.4.1 Discriminators
Sentence structure is important for constructing a complete sentence. Since the real seeds are made from perfect sentences, they maintain structural representations of sentences. Therefore, the fake seeds should capture the structural features of the real seeds. We use Bidirectional Encoder Representations from Transformer (BERT) [6] $d_{\alpha}(\cdot)$, which called Seed Structure Discriminator (SSD), to capture the structural features of sentences, and the first token is used to predict whether the seed is real or fake.
The order of tokens is also important for constructing sentences. It is possible to predict whether a sentence order representation of a seed is correct because both real and fake seeds have a dimension of (sequence length * embedding dimension). To do this, we use Bidirectional LSTM $d_{\beta}$, which called Seed Order Discriminator (SOD), to consider both forward and backward directions of sentences. Concatenated the first and the last hidden states are used to predict whether the seed is real or fake. Finally, SSD and SOD updates both the discriminator and the generator during adversarial training and their loss function can be defined as the following equation:
$\small\begin{gathered}\mathcal{L}_{D}=-\frac{1}{N}\Sigma_{i=1}^{N}\big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}\\ x=predicted,y=target\end{gathered}$ (7)
The detailed specification and figure of the discriminators can be found in Appendix A.
3.4.2 Generator Helpers
Since capturing the distribution of text embedding space is important, we conduct Seed Distribution Prediction (SDP). However, since the text embedding space cannot be directly used as a probability distribution, the seed interpretation model output is used. Specifically, if a seed passes through the frozen seed interpretation model, the output dimension of (sequence length * vocab size) is obtained by the softmax function which can be used as a probability distribution. As shown in Figure of Appendix A, the loss is calculated by Kullback-Leibler divergence with distributions of the real and the fake seeds. SDP is used only for updating the generator during adversarial training:
$\small\mathcal{L}_{SDP}=\sigma\big{(}f_{\theta}(H_{real})\big{)}\log{\frac{\sigma\big{(}f_{\theta}(H_{real})\big{)}}{\sigma\big{(}f_{\theta}(H_{fake})\big{)}}}$ (8)
where the $\sigma$ and $f_{\theta}(\cdot)$ mean softmax function and the frozen seed interpretation model.
The sentences used as the seeds are composed of tokens explained in Equation 2. Since the structures of seeds are somewhat formalized, we conduct Seed Frame Prediction (SFP). Therefore, we calculate the Mean Absolute Error (MAE) and Mean Squared Error (MSE) to make a form of a fake seed similar to a real one. If we train the fake seeds through the MAE and the MSE, the fake seeds from the generator can be blurred. However, the loss of SFP is quite small compared to that of SSD, SOD, and SDP; therefore, SFP does not harm the generator. SFP is used only for updating the generator during adversarial training:
$\small\begin{gathered}\mathcal{L}_{SFP}=\|\mu_{r}-\mu_{f}\|_{2}^{2}+\|H_{real}-H_{fake}\|_{1}\\ \mu_{r}=avg(H_{real}),\mu_{f}=avg(H_{fake})\end{gathered}$ (9)
4 Text Synthesis Experiments
4.1 Dataset
Table 1: Statistics of DailyDialog
Data Train Validation Test
Dialogues 11,118 1,000 1,000
Total sentences 87,170 8,069 7,740
Avg. sentences 7.84 8.07 7.74
TESGAN synthesizes text based on the seed obtained from the prior sentence. Therefore, the seed interpretation model should be trained with multi-turn sentence before adversarial training. In this experiment, we use DailyDialog [^1] http://yanran.li/dailydialog, data are submitted in the supplementary. [17] multi-turn conversation data, which is a dataset for training an open-domain dialogue generation model. DailyDialog consists of chit-chat-style multi-turn English conversations, and we select this data for domain-independent text synthesis. Statistics of DailyDialog are shown in Table 1.
4.2 Training Steps
Algorithm 1 Text Embedding Space GAN
Require: Seed interpretation model $f_{\theta}$; Generator $g_{\phi}$; BERT discriminator $d_{\alpha}$; LSTM discriminator $d_{\beta}$; Multi-turn data $D=\{S_{1:N}\}$; Sentence data $S_{i}=\{w^{i}_{1:L}\}$.
1:  Pre-train $f_{\theta}$ using $D$.
[…]
3:  Freeze the $f_{\theta}$.
4:  while TESGAN converges do
5:     for d-steps (during odd epoch) do
6:        Get real data from $f_{\theta}$ using $S$ with positive label 1.
7:        Make fake data from $g_{\phi}$ with negative label 0.
8:        Update $\alpha$ and $\beta$ via results of $d_{\alpha}$ and $d_{\beta}$.
9:     end for
10:     for g-steps do
11:        Make fake data from $g_{\phi}$ with positive label 1.
12:        Calculate SDP and SFP.
13:        Update $\phi$ via results of $d_{\alpha}$, $d_{\beta}$, SDP and SFP.
14:     end for
15:  end while
First, the seed interpretation model must be pre-trained by multi-turn conversation data $D=S_{1},...,S_{N}$ to interpret the seeds. As shown in Table 1, training is conducted with 11k multi-turn dialogue data, and the model achieving the highest BLEU-4 score in the validation set is selected. After pre-training the seed interpretation model, the generator and the discriminator learn through adversarial training. For adversarial training, real and fake seeds are made by the embedding part of the frozen seed interpretation model and the generator, respectively. Since the real seeds can be created from any sentence, all 87k sentences in the training set are used to make the real seeds via Equation 2. We also generate the same number of fake seeds as the real seeds to conduct adversarial training:
$\small\begin{gathered}\mathcal{D}_{\mathcal{L}}=\max_{\alpha,\beta}{\mathbb{E}_{x\sim H_{real}}\Big{[}\log{d_{\alpha,\beta}(x)}\Big{]}}\\ \mathcal{G}_{\mathcal{L}}=\max_{\phi}{\mathbb{E}_{z}\Big{[}\log{d_{\alpha,\beta}(g_{\phi}(z))}\Big{]}}+\mathcal{L}_{SDP}+\mathcal{L}_{SFP}\end{gathered}$ (10)
where $d_{\alpha,\beta}$ means SSD, SOD respectively. The sentences used as the seeds are composed of tokens explained in Equation 2. Detailed explanation of above pseudocode, optimization methods, and hyperparameters are explained in Appendix B.
4.3 Evaluation Metric
Target-oriented evaluation metrics, such as BLEU and ROUGE, are not suitable for evaluating synthetic text. This is because each synthesized sentence from random noise has no corresponding target and the generative models aim to synthesize plausible data based on real data distribution without copying the training data. Several methods were suggested that can be used in this case [20, 4].
4.3.1 Fréchet BERT Distance (FBD)
de Masson d’Autume et al. (2019) [20] proposed Fréchet Embedding Distance (FED) to evaluate the quality of synthetic text inspired by Fréchet Inception Distance (FID) [10]. Alihosseini et al. (2019) [1] proposed FBD, an improved version of FED, to measure the quality and diversity of synthesized text using pre-trained BERT. The features of real and synthesized text obtained by the pre-trained BERT are assumed to have Gaussian distributions and FBD is the distance between them:
$\small FBD=\sqrt{\|\mu_{r}-\mu_{f}\|_{2}^{2}+tr\big{(}\Sigma_{r}+\Sigma_{f}-2(\Sigma_{r}\Sigma_{f})^{0.5}\big{)}}$ (11)
where $\mu$ and $\Sigma$ show the mean vectors and the covariance matrices of the real and fake seed features.
4.3.2 Multi-Sets-Jaccard (MSJ)
Each synthesized sentence has no corresponding target; thus, we select MSJ that calculates the score between real and synthesized text sets. The Jaccard Index determines the similarity of two sets, calculating the ratio of the cardinality of their intersection to that of their union. Inspired by Jaccard Index, MSJ focuses on the similarity of the n-grams frequencies of text in the two sets of $s_{r}$ and $s_{f}$ which are real and synthesized text sets respectively:
$\small MSJ_{n}=\frac{\Sigma_{g\in G_{n}}\min\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}{\Sigma_{g\in G_{n}}\max\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}$ (12)
where $G_{n}$ and $C_{n}(g,s)$ shows the n-grams in $s_{r}\cup s_{f}$ and the normalized counts of the n-gram in set $s$. Thus MSJ can consider both quality and diversity of synthetic results.
4.3.3 Language Model score (LM)
de Masson d’Autume et al. (2019) and Caccia et al. (2020) [20, 4] proposed LM which can evaluate the quality of generated samples using a well-trained language model. LM measures generated sample quality which means that scores of the bad samples are poor under a well-trained language model. We select pre-trained GPT-$2$https://huggingface.co/docs/transformers/model_doc/gpt2 as a well-trained language model. LM is calculated as the cross-entropy value of the output and input of GPT-2.
4.3.4 Data Synthesis Ratio (DSR)
DSR considers not only data memorization ratio between training and synthesized data but also synthetic diversity itself. Short sentences identical to training data, such as "I’m fine", can be synthesized by coincidence. Therefore, sentences longer than two-thirds of the maximum sentence length that perfectly reproduce the training data are considered memorized data. Considering these conditions, we can calculate DSR using the following equation:
$\small\begin{gathered}R_{syn}=\frac{|S_{syn}-S_{train}|}{|S_{syn}|},R_{unq}=\frac{|S_{unq}|}{|S_{syn}|}\\ DSR=\frac{2*R_{syn}*R_{unq}}{R_{syn}+R_{unq}}\\ \end{gathered}$ (13)
where $S_{syn}$ and $S_{train}$ indicate synthesized and training text set respectively. $S_{unq}$ means the set of unique sentences of synthesized text results. If the synthesized sentence in $S_{syn}$ do not reproduce any of the sentences in $S_{train}$, $R_{syn}$ would be 1. Similarly, if the synthesized text in $S_{syn}$ is all unique, $R_{unq}$ will be 1. The final DSR is calculated as the harmonic mean of $R_{syn}$ and $R_{unq}$ ratio.
4.4 Baselines
Here, we compare our two models with the following approaches: Maximum Likelihood Estimation (MLE), PG-BLEU, SeqGAN, RankGAN, and MaliGAN. MLE is the pre-training result of the generator, which all the other models do before adversarial training. Since MLE is trained without adversarial training, it is shown as a baseline in Figure 2. The pre-trained generators with the lowest loss in the validation set were chosen for each method including MLE. We compare these models with our original TESGAN and P-TESGAN trained by adding zero-centered normal distribution $z$ to the generator output. We also present the results of using Gaussian random noise as input of the seed interpretation model to prove that the results of the TESGAN-based model are due to seeds created by the generator, not the seed interpretation model.
5 Results
(a) FBD $\downarrow$
(b) MS-Jaccard2 $\uparrow$
(c) MS-Jaccard3 $\uparrow$
(d) MS-Jaccard4 $\uparrow$
(e) MS-Jaccard5 $\uparrow$
(f) DSR $\uparrow$
Figure 2: Illustration showing the results of the text-GAN models. In the previous research, adversarial training is conducted after the generator pre-training. MLE is represented as a baseline because it is a supervised pre-trained generator without adversarial training.
Table 2: Performance of the models. P-TESGAN denotes the perturbed TESGAN. Numbers in parentheses indicate the training epoch of the selected model. $\dagger$ is the result of directly entering Gaussian random noise as an input to the seed interpretation model. $\ast$ denotes a metric not considered when selecting the best model. - denotes that the confidence of the result is low because the quality of the synthesized sentence is poor.
Method FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR ($R_{syn},R_{unq}$) normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
TESGAN (17) 2.899 0.150 0.079 0.042 0.021 0.967 (1, 0.936) 4.236
P-TESGAN (10) 2.274 0.131 0.066 0.032 0.014 0.841 (0.997, 0.727) 3.642
SeqGAN (1) 6.153 0.185 0.091 0.040 0.015 0.880 (0.883, 0.877) 5.094
RankGAN (1) 6.409 0.189 0.096 0.048 0.023 0.890 (0.895, 0.886) 5.123
MaliGAN (1) 21.436 0.015 0.006 0.003 0 0.030 (1, 0.015) -
PG-BLEU (1) 9.002 0.071 0.036 0.015 0.006 0.569 (0.555, 0.584) 4.584
MLE 6.284 0.193 0.096 0.045 0.021 0.955 (0.925, 0.987) 5.168
Random Noise $\dagger$ 14.142 0.080 0.036 0.016 0.006 0.930 (1, 0.869) 4.562
Table 3: Example of the synthesized sentences. P-TESGAN denotes the perturbed TESGAN.
TESGAN (17-epoch) P-TESGAN (10-epoch) Random Noise
I’m so glad you finally got on the train. Hello, Mr. Smith. I’m Mary. Anything I have called three weeks
I just lost my job. I just want to tell you the truth. Is Is Is Is Is Is Is
Yeah. You mean the network connection? It’s the end of the world. Left and go to go to go to go
What happened? What do you want to do in this company? Mr Moon, Mr Moon… Mr Moon…
So you have to wait for a while. He just broke up with Ann. are you have finished 6 items?
We use FBD, MSJ, LM, and DSR for comparing the results of the synthesized text at every epoch of adversarial training. Since FBD and MSJ require a real text set, the test set is used as the real text set. DSR is calculated with the training set as the data memorization ratio needs to be computed. FBDs of the TESGAN-based models are lower than MLE, while baselines increase after having the lowesIn terms of MSJ, as shown in Figure 1(b)-1(e), the previous studies report lower values than the TESGAN-based models at the end of adversarial training, despite having higher results in the beginning. On the other hand, MSJ results of the TESGAN-based models slightly increase during adversarial training. Moreover, some MSJ5 results of the original TESGAN are higher than MLE during adversarial training as shown in Figure 1(e). In the case of DSR shown in Figure 1(f), DSR results of the original TESGAN increase during adversarial training, and some results are higher than MLE. On the other hand, the results of the previous studies decrease during adversarial training, resulting in lower values than the TESGAN-based models in the end. As adversarial learning progresses, the results of the baselines deteriorate because the LSTM generator tends to generate only a few unique sentences.
We choose the best model of each method considering the overall results shown in Figure 2. As shown in Table 2, TESGAN and P-TESGAN at the end of 17 and 10-epoch of adversarial training are selected, respectively. For all the baselines, the 1-epoch models are selected immediately after supervised pre-training. After selecting the best models, we calculate the LM based on the text generated by each model, and the TESGAN-based models show the best results. The results of Gaussian random noise demonstrate that the TESGAN results are attributed to the seeds from the generator. Table 3 shows the synthetic text by TESGAN and Gaussian random noise.
In conclusion, MLE is a supervised pre-trained generator applied before adversarial training, but most of the result curves of the prior methods show lower performance than MLE during adversarial training. On the other hand, our TESGAN-based models have shown better results than MLE or have improved performance during adversarial training.
Error Analysis
(a) FBD $\downarrow$
(b) MS-Jaccard2 $\uparrow$
(c) MS-Jaccard3 $\uparrow$
(d) MS-Jaccard4 $\uparrow$
(e) MS-Jaccard5 $\uparrow$
(f) DSR $\uparrow$
Figure 3: Illustration of TESGAN training results. TESGANs show similar trends for every trial and SEMs decrease during adversarial training.
We also conduct three additional TESGAN training without setting a manual seed in the code to confirm the reproducibility. As a result, the overall tendency of training outcomes during adversarial training is similar. We compute the mean and the SEM of the four results and the SEM is calculated by $\frac{\sigma}{\sqrt{n}}$ (n denotes the number of experiments). Also, the SEM of each epoch decreases during adversarial training, which means that each TESGAN converges. Figure 3 shows the results of four experiments, mean, and SEM.
6 Discussion
In this experiment, we find that the performance of TESGAN depends on hyperparameters and training strategies. In the generator, the ReLU and the hyperbolic tangent (tanh) are worse than the Leaky ReLU [19]. Furthermore, deep convolutional layers and batch normalization tend to synthesize monotonous text. Thus, layers of the generator are constructed not deeply but widely and do not have batch normalization. Further discussion on the generator is in Appendix A.
It is important to find appropriate steps of the models because the discriminator cannot follow the learning of the generator if the number of generator-step (g-step) is too large compared to the number of discriminator-step (d-step) and vice versa. Commonly, since the convergence of the generator is slower than that of the discriminator, the g-step and the d-step are set to two and one, respectively. The entire pseudocode and optimization study are explained in Appendix B.
Ablation Study
Table 4: Results of the ablation study. Numbers in parentheses indicate the training epoch of the selected model. $\ast$ denotes a metric not considered when selecting the best model. - denotes that the confidence of the result is low because the quality of the synthesized sentence is poor.
Method FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
TESGAN w/o SSD (13) 2.8346 0.1377 0.0744 0.0394 0.0204 0.9497 4.1833
TESGAN w/o SOD (2) 4.7724 0.1125 0.0596 0.03115 0.0164 0.8656 4.7011
TESGAN w/o SDP (13) 38.1301 0.0580 0.0252 0.0099 0.0041 0.7390 -
TESGAN w/o SFP (16) 2.9202 0.1477 0.0790 0.0422 0.0209 0.9463 4.2339
TESGAN (17) 2.8994 0.1496 0.0789 0.0422 0.0214 0.9669 4.2361
We experiment to see the effect of the four objective functions in Section 3.4 and the results are shown in Table 4. When SOD and SDP are not used, there is a big difference in the results, indicating that the SOD and SDP are important for high-quality text synthesis. Since MSJ evaluates text based on the n-grams of tokens, the order of the synthesized text is important. Accordingly, the MSJ results of the w/o SOD are worse than those of the w/o SSD, which proves that SOD can capture the token order representations. Lastly, the results are generally high when all four objective functions are used.
Activation Function Study
Table 5: Performance according to the activation functions of the generator. $\ast$ denotes a metric not considered when selecting the best model. - denotes that the confidence of the result is low because the quality of the synthesized sentence is poor.
Activation FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
TESGAN None 47.261 0.108 0.060 0.032 0.016 0.982 -
Tanh 9.780 0.110 0.057 0.030 0.015 0.871 5.402
Sigmoid 2.899 0.150 0.079 0.042 0.021 0.967 4.236
P-TESGAN None 54.002 0.111 0.059 0.030 0.015 0.958 -
Tanh 20.158 0.118 0.061 0.031 0.015 0.937 -
Sigmoid 2.274 0.131 0.066 0.032 0.014 0.841 3.642
Table 6: Synthesized sentences by tanh and non-use cases in Table 5. P-TESGAN denotes the perturbed TESGAN.
TESGAN with Tanh P-TESGAN with Tanh
You are a little You ’ re a book?
I ’ m sorry to see you off. You ’ Ve come You are late.
I ’ m sorry. I ’ m doing ’ t “ all day ’ s
You ’ d like a tour to see the dentist. I don’t know what time it is?
You are late. I ’ m sorry to hear this!
TESGAN without activation P-TESGAN without activation
I ’ d like to say it! I ’ d like to I ’ s a big, that ’ s right.
Yes, do you want to buy? I like the back ones. They look like a shop.
I ’ s right over there? I ’, this ’, this ’! be real,
What’s the matter? I have a problem with my English textbooks.
I got a bite the food? I ’ s faster, George. I ’ d like to go
The results are different depending on the activation functions used at the end of the seed making process. We conduct experiments on sigmoid, tanh, and non-use cases during the seed making process, and their results are shown in Table 5. Table 6 shows the quality of the synthetic text for tanh and non-use cases, and the results are worse than those using sigmoid in Table 3. However, according to Table 5, DSR results of non-use case is higher than the sigmoid case. Thus, we can see that a higher DSR does not always mean good quality because DSR only considers data memorization. Therefore, we select the model using the sigmoid activation function, which has better results for FBD and MSJ, and moderately high DSR.
Case Study
Figure 4: DSR results according to the number of seed interpretation model parameters.
Pre-trained GPT-2 which has 124M parameters used as the seed interpretation model is a powerful language model. Therefore, we need to confirm whether the low data memorization comes from transfer learning or the TESGAN framework. We trained three smaller seed interpretation models from scratch to measure the data memorization and they have 54M, 75M, and 96M parameters each. As shown in Figure 4, DSR is high regardless of the number of model parameters during adversarial training. The results in Figure 4 indicate that the low data memorization comes from the TESGAN framework.
7 Conclusion
Herein, we proposed a novel unsupervised text synthesis framework, TESGAN. TESGAN enabled the gradient backpropagation of natural language discrete tokens by creating a continuous text embedding space called a seed. In most text-GAN studies, data memorization had been inevitable because the generator have to be pre-trained with an autoregressive way before adversarial training. Therefore, we suggested TESGAN, which removed the data memorization issue by applying an unsupervised GAN framework that does not directly refer to the training data. TESGAN improved text synthesis performance during adversarial training and resulted in the best FBD and LM. Also, TESGAN showed the lowest data memorization ratio and the case study proved that these results are attributed to the TESGAN framework. Furethermore, ablation study shows the importance of the four objective functions. This paper implies the potential of continuous embedding spaces as well as discrete tokens for text synthesis with unsupervised learning. By combining the perspective of viewing text as a continuous space with publicly available Large Language Models [25], models will synthesize more expressive sentences, and we hope many follow-up studies will come out.
Appendix 0.A Details of Models
(a) Generator
(b) Seed Structure Discriminator (SSD)
(c) Seed Order Discriminator (SOD)
(d) Seed Distribution Prediction (SDP)
Figure 5: Illustrations of the generator, two discriminators, and SDP. In the generator, P-TESGAN makes perturbed seeds by adding zero-centered normal distribution noise $z$ (gray) to the output (blue) from the generator. SSD predicts whether the seed is real or fake using the $[CLS]$ special token. SOD considers both forward and backward contexts of the seed. SDP is used to improve the fake seeds of the generator during adversarial training by reducing the distance between real and fake seed distributions.
0.A.1 Seed Interpretation Model
The seed interpretation model $f_{\theta}(\cdot)$ is necessary to predict subsequent sentences a seed makes. Therefore, the seed interpretation model must be trained with multi-turn sentences in an autoregressive way Our interpretation model inherits the 12-layer GPT-2, derived from the decoder of the transformer language model [26], and has 124M parameters. We used the model achieving the highest NLTK BLEU-4 in the validation set in Table 1 as the seed interpretation model.
0.A.2 Generator
The generator $g_{\phi}(\cdot)$ consists of two 1D transposed convolutional layers and has 3.3M parameters. The first and the second layers conduct convolution with 128 and 16 filters, respectively. Since sentences vary according to types of tokens and their order, forms of the real seed $H_{real}$ are also varied. The generator generates seeds from the uniform distribution noise $X$ with an interval of $[-10,10)$ to make diverse forms of the seeds. Furthermore, the fake seeds, results of deep convolutional layers and batch normalization, tend to synthesize only monotonous sentences. Thus, layers of the generator are constructed not deeply but widely and the generator does not have batch normalization layers.
Leaky ReLU is used as the activation function between the two convolutional layers. When the ReLU is used, Dying ReLU occurs; the negative values become zero, making it unsuitable for text synthesis where diversity is important. Also, tanh is not adequate because of the gradient vanishing problem. Figure 4(a) shows convolutional processes of the generator.
0.A.3 Seed Structure Discriminator (SSD)
Sentence structure is important for constructing a complete sentence. For example, “I love you so much” is structurally error-free, but “I love like so much” and “I love” are not. Because real seeds are created from perfect sentences, they retain the structural representation of sentences. Therefore it is important that the fake seeds should capture the structural features of the real seeds.
We assume that every sentence can be the first sentence in multi-turn cases. Thus, the real seeds are obtained from sentences where the $[CLS]$ token is inserted at the beginning. We use the 2-layer BERT $d_{\alpha}(\cdot)$ to capture the structural features of sentences, and the $[CLS]$ token is used to predict whether the seed is real (label 1) or fake (label 0). In addition, real and fake seeds do not pass through the embedding part of the BERT because they are already embedding spaces. Finally, the BERT used in SSD has 54M parameters and SSD is used to update both the discriminator and the generator during adversarial training. The detailed figure of SSD is shown in Figure 4(b).
0.A.4 Seed Order Discriminator (SOD)
The order of tokens is important for constructing sentences. For example, “I love you so much” is syntactically correct, but “I you love so much” and “I love you much so” are not. To do this, we use a 2-layer Bidirectional LSTM to consider both forward and backward directions of sentences and the model has 24M parameters. Concatenated hidden states of the last token ($[SEP]$ or $[PAD]$) and the first token ($[CLS]$) are used to predict whether the seed is real (label 1) or fake (label 0). Finally, SOD updates both the discriminator and the generator during adversarial training. The detailed figure of SOD can be found in Figure 4(c).
Appendix 0.B Reproducibility
0.B.1 Optimization
The TESGAN framework performs adversarial training after training the seed interpretation model with multi-turn data. The multi-turn data for seed interpretation model training are limited to a maximum of four turns, and the maximum length of each sentence is 16. In the adversarial training phase, the sentences used as the seeds are composed of tokens explained in Equation 2, and the length of each sentence is set to 16 including special tokens. The discriminators are updated by the loss of SSD and SOD during adversarial training. Also, the generator is updated not only by the loss of SSD and SOD but also that of SDP and SFP.
Additionally, we find that the convergence of the model depends on the parameter update rate of the discriminators and the generator. According to Algorithm 1, the discriminators update parameters only at odd training epochs so that the generator can follow the learning of the discriminators; because the convergence of the generator is slower than that of the discriminator. If the discriminators are updated every training epoch in the same way as the generator, adversarial training is unbalanced.
Also, we conduct additional experiments by changing the number of the generator updates from one to three for each mini-batch step. If the generator updates only once per step as same as the discriminators, the generator cannot follow the learning of the discriminators. On the other hand, if the generator updates three times per step, the discriminators cannot follow the learning of the generator. Therefore, we update the generator twice per step. Consequently, the generator is updated four times more than the discriminators per two epochs. Algorithm 1 shows the entire training process of TESGAN.
0.B.2 Hyperparameters
The 12-layer GPT-2 is used as the seed interpretation model, and both hidden and embedded dimensions are 768. We adopt the byte-pair-encodings (BPE) [24] tokenizer with 50,260 vocabularies in the seed interpretation model. We suggest using the Adam optimizer [14] with 0.001 learning rate, and 100 mini-batch sizes for training the seed interpretation model.
The fake seeds are generated from the uniform distribution noise $X$ with an interval of $[-10,10)$ by the generator, which has two convolutional processes. In addition, the Leaky ReLU with slope 0.5 and the sigmoid are used in the middle and the end of the generator, respectively. We suggest the Adam optimizer with 0.0002 learning rate because the generator has difficulty converging when the learning rate exceeds 0.0004. The BERT and the LSTM models, used as SSD and SOD respectively, consist of two layers and their hidden dimension is 768. Both discriminators use the Adam optimizer with 0.0005 and 0.001 learning rate, respectively. When the learning rates of the discriminators are larger than the proposed values, adversarial learning is unbalanced. Also, a mini-batch size is set to 128 during adversarial training. Lastly, all the above experiments took place on a machine with Ubuntu 18.04.5 and an NVIDIA RTX 3090 GPU.